# SAC: Neural Speech Codec with Semantic-Acoustic Dual-Stream Quantization

### v1

@@ Appendix D Mel-Spectrogram Reconstructions Comparison among Codecs
To further analyze the speech disentanglement capability of different speech codecs, we visualize mel-spectrograms reconstructed under various configurations using a randomly selected speech sample from the LibriSpeech test-clean set (identical to the one used in the main paper). In particular, we additionally include SpeechTokenizer (Zhang et al., 2023) for comparison, which achieves effective semantic–acoustic disentanglement through its multi-layer residual vector quantization (RVQ) structure, by reconstructing speech from tokens of different RVQ layers.
@@
In these settings, full reconstruction refers to speech reconstructed using all discrete speech tokens. The semantic-only reconstruction denotes cases where only semantic tokens are used, while acoustic embeddings are masked with zeros, as in semantic–acoustic decoupled codecs such as SAC and SemantiCodec (Liu et al., 2024). In SpeechTokenizer, the first RVQ layer is guided by semantic distillation to primarily encode linguistic information; therefore, reconstruction from the first-layer tokens (denoted as RVQ-1) can be considered analogous to semantic-only reconstruction. Conversely, acoustic-only reconstruction uses only the acoustic tokens while masking semantic embeddings. In SpeechTokenizer, the modeling of acoustic details is pushed to the deeper RVQ layers (layers 2–8); thus, reconstructions based on tokens from these layers (denoted as RVQ-2:8) correspond to the disentangled acoustic component.
@@
In the semantic-only reconstruction, SAC’s harmonics almost vanish, preserving only a coarse energy contour of speech while effectively removing formant-related acoustic features. This confirms that the semantic stream in SAC is minimally influenced by the acoustic stream during decoding, achieving clean disentanglement of non-linguistic information. SemantiCodec, however, still exhibits visible harmonic and formant structures, suggesting considerable residual acoustic details. This can be attributed to its design: the semantic tokens in SemantiCodec are derived from AudioMAE (Huang et al., 2022), which is trained on general audio rather than speech data, making it less capable of separating semantic content from acoustic cues. As a result, its semantic tokens encode general informative features rather than purely linguistic representations. SpeechTokenizer’s RVQ-1 reconstruction also shows weakened low-frequency harmonics but retains high-frequency spectral density, indicating that the first RVQ layer still carries some acoustic features.

### v2

@@ Appendix I Mel-Spectrogram Reconstructions Comparison among Codecs
To further analyze the speech disentanglement capability of different speech codecs, we visualize mel-spectrograms reconstructed under various configurations using a randomly selected speech sample from the LibriSpeech test-clean set (identical to the one used in the main paper). In particular, we additionally include SpeechTokenizer (speechtokenizer) for comparison, which achieves effective semantic–acoustic disentanglement through its multi-layer residual vector quantization (RVQ) structure, by reconstructing speech from tokens of different RVQ layers.
@@
In these settings, full reconstruction refers to speech reconstructed using all discrete speech tokens. The semantic-only reconstruction denotes cases where only semantic tokens are used, while acoustic embeddings are masked with zeros, as in semantic–acoustic decoupled codecs such as SAC and SemantiCodec (semanticodec). In SpeechTokenizer, the first RVQ layer is guided by semantic distillation to primarily encode linguistic information; therefore, reconstruction from the first-layer tokens (denoted as RVQ-1) can be considered analogous to semantic-only reconstruction. Conversely, acoustic-only reconstruction uses only the acoustic tokens while masking semantic embeddings. In SpeechTokenizer, the modeling of acoustic details is pushed to the deeper RVQ layers (layers 2–8); thus, reconstructions based on tokens from these layers (denoted as RVQ-2:8) correspond to the disentangled acoustic component.
@@
In the semantic-only reconstruction, SAC’s harmonics almost vanish, preserving only a coarse energy contour of speech while effectively removing formant-related acoustic features. This confirms that the semantic stream in SAC is minimally influenced by the acoustic stream during decoding, achieving clean disentanglement of non-linguistic information. SemantiCodec, however, still exhibits visible harmonic and formant structures, suggesting considerable residual acoustic details. This can be attributed to its design: the semantic tokens in SemantiCodec are derived from AudioMAE (audiomae), which is trained on general audio rather than speech data, making it less capable of separating semantic content from acoustic cues. As a result, its semantic tokens encode general informative features rather than purely linguistic representations. SpeechTokenizer’s RVQ-1 reconstruction also shows weakened low-frequency harmonics but retains high-frequency spectral density, indicating that the first RVQ layer still carries some acoustic features.